# Reading Images Like Texts: Sequential Image Understanding in Vision-Language Models

### v1

@@ 5.1 Token compression based on the token map
We test our method on LLaVA-1.5-7B and Qwen2.5-VL-7B. For the training of the visual decoder, we use the images from the train set of GQA and TextVQA (Singh et al., 2019), containing 148k and 22k samples, respectively. For evaluation, we use GQA, TextVQA, VQA (Antol et al., 2015) and ScienceQA (Lu et al., 2022) for visual question answering. We also use MMBench (Liu et al., 2024b) and POPE for the test of the general capabilities and hallucination. Details for training and evaluation are presented in Appendix A.6. As shown in Table 1, we implement three methods: (1) method 1 is the original algorithm; (2) method 2 deletes all the visual embeddings that are decoded into punctuation marks with a highest probability among all tokens in the vocabulary; (3) method 3 is based on method 2, while it only remove those with top-2 text tokens both being punctuation, which are more likely to be meaningless visual tokens. Results show that our method could not only reduce sequence length for faster inference, but also control performance loss within an acceptable range. Details of this are shown in Algorithm 2. Further discussions are presented in Appendix A.6.
@@ 5.2 Enhancing spatial reasoning with rectified RoPE
According to the discussions in Section 4.2, the component for positional discrimination (e.g., $\pm[(q_{A}^{(0)}k_{B}^{(1)}-q_{A}^{(1)}k_{B}^{(0)})sin(m\theta)]v_{B}$) has a relatively small magnitude. To verify this, we split the dot-product attention into two parts: $q\cdot k^{\top}=(q^{X},q^{Y})(k^{X},k^{Y})^{\top}=q^{X}(k^{X})^{\top}+q^{Y}(k^{Y})^{\top}$, and compute the contribution to attention scores from the two axes by concatenating the two parts and applying Softmax to them together. For the $X$-axis and $Y$-axis, we separately select their “attention scores” from the “satellite” to the “nucleus” and average them across all attention heads and tokens (Equation 53, 54). The test is conducted on What’s Up B and Qwen2-VL-2B. Results in Figure 24 show that for ”left” and ”right” (in the $X$-axis), attention from $Y$-axis is about 1.5 times larger than the X-axis, while the latter contains the key information. On the other hand, as the RoPE frequency $\theta_{i}=b^{-2i/d}$ decays rapidly when the dimension group index $i$ increases, the dimensions corresponding to large values of $i$ become highly insensitive to positional changes. For example, when $i=\frac{d}{2}$ and RoPE base $b=10000$ (in Qwen2-VL), for relative distances of -50 and 50, the difference in their sine function values, $2sin(50\theta_{d/2})$, is merely around 0.01. To address this, we propose RoPE scaling, which adaptively scales the relative distances via a modification: $(\theta_{i})^{\prime}=\theta_{i}\cdot g(i)$, where
@@
In Equation 7, $d$ is the dimension of an attention head, $\alpha$ serves as the scaling magnitude, while $p$ is used to control the scaling factor $g(i)$ such that it applies minimal scaling for small $i$ and significant scaling for large $i$ (see Figure 25(b) for illustration). This compensates for the loss of positional information caused by frequency decay. We test our method (training-free / fine-tune on 60k samples from the GQA training set) on the Qwen2-VL series of models and evaluate on What’s Up, VSR (Liu et al., 2023a), and COCO/GQA-spatial (Kamath et al., 2023). The selection of $\alpha$ and $p$, along with the training and evaluation setups, are presented in Appendix A.7. Results in Table 2 show that RoPE scaling achieves competitive performance across multiple benchmarks. We also test our method on MMBench and find that the model performance in general capabilities is even better with RoPE scaling, demonstrating its potential as a trick to enhance spatial reasoning ability.
6 Related works
Interpretability of VLMs A line of work aims to interpret model components. Gandelsman et al. (2023; 2024) find the neurons and attention heads in CLIP that correspond to real-world concepts. Hugofry (2024); Daujotas (2024); Rao et al. (2024) extract CLIP features using sparse autoencoders. Neo et al. (2024); Sonia Joseph (2024) interpret activations via logit lens. Another line of work employs the idea of causal tracing. Rajaram et al. (2024); Palit et al. (2023); Basu et al. (2024) investigate information storage and transfer in CLIP, BLIP and other VLMs via visual circuits. Recently, Neo et al. (2024); Zhang et al. (2024b; 2025) use blocking-based interventions to study the information flow in VLMs. Other approaches provide visualizations based on gradients (e.g. (Zhang et al., 2024b)), attention (Ben Melech Stan et al., 2024) or causal inference (Li et al., 2025).
@@ 7 Conclusion
We investigated image understanding in vision-language models from the perspectives of object recognition (the “what” way) and spatial perception (the “where” way), and applied our analytical findings to practice. Our work offers new insights for both better understanding the mechanisms of visual information processing in VLMs and designing superior algorithms or model architectures. Limitations of this work are detailed in Appendix A.1.
@@ 8 Reproducibility Statement
To ensure reproducibility of this work, key supporting materials are distributed in the main text, appendix, and supplements. The anonymous source code for our proposed model/algorithm is available in Supplementary Material. For theoretical analysis for spatial perception, key assumptions and results are presented in Section 4 of the main text, with derivation process detailed in Appendix A.5.1. Dataset details—including sources, preprocessing, and splitting—are overviewed in the main text and detailed in the Appendix and Supplementary Material. These materials enable accurate replication of our results. Finally, LLMs are used to aid our writing of this paper.
Appendix A Appendix
A.1 Limitations
@@
A.2 Details for the Rotary Position Embedding
A.2.1 1D RoPE in language models
@@ A.2 Details for the Rotary Position Embedding
A.2.2 2D RoPE in vision-language models
@@
A.2.3 Engineering details of RoPE
@@ A.3 Details for the models
A.4 Analysis based on logit lens
A.4.1 Analysis on the representation similarity
@@
A.4.2 Details about the token map
@@
A.4.3 Details about the segmentation map
@@
A.4.4 Further discussions about the Gestalt Recognition
@@
A.5 Details of the analysis on spatial perception
A.5.1 Theoretical analysis
@@
A.5.2 Empirical analysis on the representation of a single object
@@
A.5.3 Empirical analysis on the direction vectors
In Section 4.2, we formally defined the direction vector; here, we detail its calculation method. Contrary to our theoretical assumption, in a real image, an object does not correspond to just one image patch. As shown in Figure 7(a), a single object can comprise tens to over a hundred image patches. In this case, we define the direction vector as $v^{r}=h_{o_{S}}^{r}-h_{o_{N}}^{r}=\operatorname{Mean}(V_{S}^{r})-\operatorname{Mean}(V_{N}^{r})(r\in\mathcal{R})$, where $\operatorname{Mean}$ represents mean pooling, and the definitions of $V_{S}^{r}$ and $V_{N}^{r}$ are detailed in Appendix A.5.2. In our actual calculations, to better validate our theory and visualize the direction vectors, we did not fix the representation $h_{o}^{r}$ of an object o to a single deterministic value $\operatorname{Mean}(V_{o}^{r})$. Instead, we randomly sample $n\;(n<|V_{o}^{r}|)$ patch embeddings from $V_{o}^{r}$ and then perform mean pooling.
@@ A.5.3 Empirical analysis on the direction vectors
A.6 The token compression algorithm based on run-length encoding (RLE)
In Section 5.1, we propose a token compression method based on Run-Length Encoding (RLE). Prior methods primarily include those based on the similarity between visual tokens (Wen et al., 2025b; Alvar et al., 2025; Jeddi et al., 2025) and those based on instruction relevance (Chen et al., 2024a; Zhang et al., 2024c; Xing et al., 2024). Similarity-based approaches compress visual tokens according to their similarity. Such methods depend on the chosen similarity metric, and commonly used metrics like cosine similarity are not universally applicable. As discussed in Appendix A.4.1, in the deeper layers of a ViT, the representations of visual tokens often contain information from other tokens, and this entanglement becomes increasingly complex with depth. This makes it difficult to compute the similarity between visual tokens using simple metrics in a linear space.

### v2

@@ 5.1 Token compression based on the token map
We test our method on LLaVA-1.5-7B and Qwen2.5-VL-7B. For the training of the visual decoder, we use the images from the train set of GQA and TextVQA (Singh et al., 2019), containing 148k and 22k samples, respectively. Training such a visual decoder is quite simple. As detailed in Appendix G, it takes approximately 5 hours on an A40 GPU and is highly data efficient. For LLaVA-1.5-7B, its visual decoder only requires training for $6k$ steps with a batch size of $16$, using less than $100k$ of unlabeled data. For evaluation, we use GQA, TextVQA, VQA (Antol et al., 2015) and ScienceQA (Lu et al., 2022) for visual question answering. We also use MMBench (Liu et al., 2024b) and POPE for the test of the general capabilities and hallucination. As shown in Table 1, we implement three methods: (1) method 1 is the original algorithm; (2) method 2 deletes all the visual embeddings that are decoded into punctuation marks with a highest probability among all tokens in the vocabulary; (3) method 3 is based on method 2, while it only remove those with top-2 text tokens both being punctuation, which are more likely to be meaningless visual tokens. Details of these choices are shown in Algorithm 2. Further details of training and evaluation are presented in Appendix G.
Results show that our method could not only reduce sequence length for faster inference, but also control performance loss within an acceptable range. Compared with previous methods that are based on instruction relevance (Chen et al., 2024a; Zhang et al., 2024c; Lin et al., 2025; Zhu et al., 2024), though they seem to be better in terms of compression ratio and task performance, they are impractical in real-world scenarios because they cannot use FlashAttention (Dao et al., 2022) and require re-computing the relevance to the user prompt for every inference. Our method is more similar to the methods that focus solely on the image itself. The most representative is ToME (Bolya et al., 2022). While they measure the similarity between visual tokens via the keys from the ViT’s attention, our method is more refined because it converts an image to natural language and merges similar tokens directly via text strings. Thus it has a lower compression ratio than ToME, but better performance on downstream tasks. More comparisons are provided in Appendix G.
@@ 5.2 Enhancing spatial reasoning with rectified RoPE
According to the discussions in Section 4.2, the component for positional discrimination (e.g., $\pm[(q_{A}^{(0)}k_{B}^{(1)}-q_{A}^{(1)}k_{B}^{(0)})sin(m\theta)]v_{B}$) has a relatively small magnitude compared with the common terms shared between the object representations of a pair of spatial relationships. This is because the values of trigonometric functions, which carry relative distance information, are always less than 1. To verify this, we focus on attention weights that RoPE directly influences. We split the dot-product attention into two parts: $q\cdot k^{\top}=(q^{X},q^{Y})(k^{X},k^{Y})^{\top}=q^{X}(k^{X})^{\top}+q^{Y}(k^{Y})^{\top}$, and compute the contribution to attention scores from the two axes by concatenating the two parts and applying Softmax to them together. For the $X$-axis and $Y$-axis, we separately select their “attention scores” from the “satellite” to the “nucleus” and average them across all attention heads and tokens (defined in Equation 53, 54). Details of this process are shown in Appendix H. The test is conducted on What’s Up B and Qwen2-VL-2B. Results in Figure 24 show that for ”left” and ”right” (in the $X$-axis), attention from $Y$-axis is about 1.5 times larger than the X-axis, while the latter contains the key information for the determination of the positional relationship between two objects.
On the other hand, as the RoPE frequency $\theta_{i}=b^{-2i/d}$ decays rapidly when the dimension group index $i$ increases, the dimensions corresponding to large values of $i$ become highly insensitive to positional changes. For example, when $i=\frac{d}{2}$ and RoPE base $b=10000$ (in Qwen2-VL), for relative distances of -50 and 50, the difference in their sine function values, $2sin(50\theta_{d/2})$, is merely around 0.01. These issues can all lead to critical information being “drowned out” by other irrelevant information when the model determines relative positional relationships between objects.
@@
To address the above issues, we propose RoPE scaling, which adaptively scales the relative distances via a modification: $(\theta_{i})^{\prime}=\theta_{i}\cdot g(i)$, where
@@
In Eq. 7, $d$ is the dimension of an attention head, $\alpha$ serves as the scaling magnitude, while $p$ is used to control the scaling factor $g(i)$ such that it applies minimal scaling for small $i$ and significant scaling for large $i$ (see Figure 25(b) for illustration). This compensates for the loss of positional information caused by frequency decay. We test our method (training-free / fine-tune on 60k samples from the GQA training set) on the Qwen2-VL series of models and evaluate on What’s Up, VSR (Liu et al., 2023a), and COCO/GQA-spatial (Kamath et al., 2023). The selection of $\alpha$ and $p$, along with the training and evaluation setups, are presented in Appendix H. Results in Table 2 show that RoPE scaling achieves competitive performance across multiple benchmarks. We also test our method on MMBench and find that the model performance in general capabilities is even better with RoPE scaling, demonstrating its potential as a trick to enhance spatial reasoning ability. However, it serves only as a localized remedy for potential limitations in RoPE. We expect developments in positional encoding that can more effectively capture relative spatial relationships in the future.
6 Related work
Interpretability of VLMs A line of work aims to interpret model components. For example, Goh et al. (2021); Gandelsman et al. (2023; 2024) find the neurons and attention heads in CLIP that correspond to real-world concepts. Elhage et al. (2022) demonstrate that neurons are polysemantic due to the existence of superposition, and sparse autoencoders (SAEs) are used to address this issue in LLMs (Bricken et al., 2023). Similarly, Hugofry (2024); Daujotas (2024); Rao et al. (2024) extract features in CLIP using SAEs. Similar to our work, Neo et al. (2024); Sonia Joseph (2024) interpret activations via logit lens. Sonia Joseph (2024) directly maps visual tokens to labels via the classification head in ViT. Neo et al. (2024) maps the visual tokens to texts via the language head. Our work is based on theirs, while we go a step further to systematically investigate VLMs layer by layer, and provide a detailed analysis of the dynamic visual information processing in a VLM.
Another line of work employs the idea of causal tracing originally developed for interpreting LLMs (Meng et al., 2022). Rajaram et al. (2024); Palit et al. (2023); Basu et al. (2024) investigate information storage and transfer in CLIP, BLIP and other VLMs via visual circuits. Recently, Neo et al. (2024); Zhang et al. (2024b; 2025) use blocking-based interventions to study the information flow in VLMs. Other approaches provide visualizations for the model internals. Some work provides heatmaps based on gradients (Selvaraju et al., 2017; Chefer et al., 2021; Zhang et al., 2024b), attention (Ben Melech Stan et al., 2024) or causal inference (Li et al., 2025).
@@ 7 Conclusion
We investigated image understanding in VLMs from the view of object recognition (the “what” way) and spatial perception (the “where” way). Based on the findings, we proposed an instruction-agnostic token compression method for inference acceleration, and RoPE scaling to improve the spatial reasoning capabilities of VLMs based on 2D RoPE. Our work offers new insights for both better understanding the mechanisms of visual information processing in VLMs and designing superior algorithms or model architectures. Limitations of this work are detailed in Appendix A.
8 Acknowledgement
We thank the anonymous reviewers for their insightful comments. The research is supported by the Natural Science Foundation of Beijing, China (Grant No. L247010).
@@ 9 Reproducibility Statement
To ensure reproducibility of this work, key supporting materials are distributed in the main text, appendix, and supplements. The anonymous source code for our proposed model/algorithm is available in Supplementary Material. For theoretical analysis for spatial perception, key assumptions and results are presented in Section 4 of the main text, with derivation process detailed in Appendix F.1. Dataset details—including sources, preprocessing, and splitting—are overviewed in the main text and detailed in the Appendix and Supplementary Material. These materials enable accurate replication of our results. Finally, LLMs are used to aid our writing of this paper.
Appendix A Limitations
@@
Appendix B The use of large language models
LLMs are used to aid our writing of this paper, primarily for checking vocabulary, correcting grammar, and polishing the prose.
Appendix C Details for the Rotary Position Embedding
C.1 1D RoPE in language models
@@ Appendix C Details for the Rotary Position Embedding
C.2 2D RoPE in vision-language models
@@
C.3 Engineering details of RoPE
@@ Appendix D Details for the models
Appendix E Analysis based on logit lens
E.1 Analysis on the representation similarity
@@
E.2 Details about the token map
@@
E.3 Details about the segmentation map
@@
E.4 Further discussions about the Gestalt Recognition
@@
Appendix F Details of the analysis on spatial perception
F.1 Theoretical analysis
@@
F.2 Empirical analysis on the representation of a single object
@@
F.3 Empirical analysis on the direction vectors
In Section 4.2, we formally defined the direction vector; here, we detail its calculation method. Contrary to our theoretical assumption, in a real image, an object does not correspond to just one image patch. As shown in Figure 7(a), a single object can comprise tens to over a hundred image patches. In this case, we define the direction vector as $v^{r}=h_{o_{S}}^{r}-h_{o_{N}}^{r}=\operatorname{Mean}(V_{S}^{r})-\operatorname{Mean}(V_{N}^{r})(r\in\mathcal{R})$, where $\operatorname{Mean}$ represents mean pooling, and the definitions of $V_{S}^{r}$ and $V_{N}^{r}$ are detailed in Appendix F.2. In our actual calculations, to better validate our theory and visualize the direction vectors, we did not fix the representation $h_{o}^{r}$ of an object o to a single deterministic value $\operatorname{Mean}(V_{o}^{r})$. Instead, we randomly sample $n\;(n<|V_{o}^{r}|)$ patch embeddings from $V_{o}^{r}$ and then perform mean pooling.
@@ F.3 Empirical analysis on the direction vectors
Appendix G The token compression algorithm based on run-length encoding (RLE)
In Section 5.1, we propose a token compression method based on Run-Length Encoding (RLE). Prior methods primarily include those based on the similarity between visual tokens (Wen et al., 2025b; Alvar et al., 2025; Jeddi et al., 2025) and those based on instruction relevance (Chen et al., 2024a; Zhang et al., 2024c; Xing et al., 2024). Similarity-based approaches compress visual tokens according to their similarity. Such methods depend on the chosen similarity metric, and commonly used metrics like cosine similarity are not universally applicable. As discussed in Appendix E.1, in the deeper layers of a ViT, the representations of visual tokens often contain information from other tokens, and this entanglement becomes increasingly complex with depth. This makes it difficult to compute the similarity between visual tokens using simple metrics in a linear space.
@@ Appendix G The token compression algorithm based on run-length encoding (RLE)
For inference efficiency, we provide here a theoretical derivation for the compression ratio of our method. Since we only add an extra visual decoder after the modality connector compared to the original model, we ignore the computational cost of the ViT and the modality connector, and only consider the LLM. Furthermore, we focus only on the main overhead—the matrix multiplications within the Attention and MLP modules—while ignoring the text embedding, Softmax, activation functions, and the final language head.
Assume the LLM has $L$ layers, the model dimension is $d$, and the intermediate dimension of the MLP is $d^{\prime}$. The visual decoder $\varphi$ has a size of $\mathbb{R}^{d\times|\mathcal{V}|}$, where $\mathcal{V}$ is the vocabulary. We consider the input of a single image, where the visual sequence length is $n$, and we ignore the length of the instruction. The computational cost of the original model is:C = (8nd^2+4n^2d+4ndd’)L  (FLOPs)When the compression ratio is $r$, the computational cost of our method is:C’=[8n(1-r)d^2+4n^2(1-r)^2d+4n(1-r)dd’]L+2nd—V—  (FLOPs)Therefore, we can measure inference efficiency by $\frac{C’}{C}$. For LLaVA-1.5-7B, the compression ratio $r$ of our method can reach $0.58$, corresponding to $\frac{C’}{C}=43.82$%, which is highly efficient.